Assume the following information:

90-day Ghana interest rate = 4%

90-day South African interest rate = 3%

90-day forward rate of South African rand = GHS 0.3500

Spot rate of South African rand = GHS 0.3550

Assume that the Osei Bonsu Co. in Ghana will need 300,000 rand in 90 days to pay for

imports from South Africa. It wishes to hedge this payables position. Would it be better

off using a forward hedge or a money market hedge? Substantiate your answer with

estimated costs for each type of hedge.


1
Expert's answer
2021-08-03T15:21:05-0400

Forward Hedge Calculation.

Forward Hedge = Spot Rate (1 + Domestic Interest Rate)/(1 + Foreign Interest Rate)

"Forward Hedge = 0.3550*(1+ 0.04 )\/(1+0.03)"

"=0.3550*(1.04\/1.03)"

"=0.3550*1.0097"

Forward Rate Hedge = R 0.3584


Money Market Hedge.

Spot Rate = GHS 0.3550

R1 = GHS 0.3550

R300,000 =

"(R300,000 * GHS0.3550)\/R1" = GHS106,500

Domestic Currency at spot rate = GHS106,500

Placing the domestic currency on deposit at prevailing interest rates.

GHS106,500 * 4% = GHS4,260

At maturity = GHS 4,260

"(4,260*1)\/0.3550"

=R12,000

Market Hedge = R12,000/R106,500

= R 0.1127


By using the money market hedge, the company would have effectively locked in a 90-day forward rate of 0.3584. Another potential reason could be that the firm is small to obtain a forward currency facility from its banker or perhaps it did not get a competitive forward rate and decided to structure a money market hedge instead.